3 1/5 + b = 6 1/10
b =

Found 2 solutions by jim_thompson5910, lynnlo:
Answer by jim_thompson5910(35256)

3 1/5 + b = 6 1/10

b = 6 1/10 - 3 1/5

b = 6 1/10 - 3 2/10

b = 5 11/10 - 3 2/10

b = (5-3) (11/10 - 2/10)

b = 2 (11-2)/10

b = 2 9/10

Answer by lynnlo(4176)